Patty and Rick deliver a total of 72 papers each morning. Patty delivers 6 more than two times as many as Rick. How many papers does Rick deliver? Show your work.
Answer by ewatrrr(24785)

 
Hi,
Patty delivers 6 more than two times as many as Rick
Let x and (2x + 6) represent #of Papers for Rick and Patty respectively
Question states***
Patty and Rick deliver a total of 72 papers each morning
x +(2x + 6) = 72
3x+6 = 72
3x = 66
x = 22, number Rick Delivers, Patty delivers 50 (2·22 + 6)
CHECKING our Answer***
22 + 50 = 72